please solve the systems of linear equations:
y= (1/2)
y= cosx

Answer by stanbon(75887)

please solve the systems of linear equations:
y= (1/2)
y= cosx
----
cos(x) = 1/2
---
x = 60 degrees or pi/3
y = 1/2
-----------------
Cheers,
Stan H.